OpenStudy (anonymous):

The ordered pairs (1,5) (2,25) (3,135) (4,625) and (5,3125) represent a function. What is a rule that represents this function? a.) y=x^5 b.)y=5x c.)y=5^x d.)y=x+5

OpenStudy (anonymous):

@itsmylife

OpenStudy (anonymous):

I think the 135 should be 125?

OpenStudy (anonymous):

Yeah, my bad. lol

OpenStudy (campbell_st):

well just substitute a few values... it can't be (d) as 5 = 1 + 5 nope after substituting the point (1, 5) and in (b) substitute (1, 5) so you get y = 5*1 so that works try (2, 25) 25 = 5*2 nope 25 doesn't equal 10 so its not b try (a) substitute (1, 5) 5 = 1^5 nope you get 5 = 1 which isn't true look at (c) does 5 = 5^1....? 25 = 5^2...? I'll let you decide

OpenStudy (anonymous):